OpenStudy (anonymous):

Write the equation of the line that passes through (-3, 4) with a slope of 2 in point-slope form.

jimthompson5910 (jim_thompson5910):

point slope form is y - y1 = m(x - x1) where m = slope (x1,y1) is the given point the line goes through

jimthompson5910 (jim_thompson5910):

hopefully that's enough to get you going let me know if otherwise